Scenario - A 65-employee firm has the following documented in the OSHA 300 record for the year:

            Case 1 - Work-related injury; no workdays lost

            Case 2 - Work-related illness; 5 days away from work

            Case 3 - Work-related injury; one workday away from work; nine more days in restricted work activity

            Case 4 - Injury; able to stay at work, but for 10 days employee was unable to do regular job and was assigned to another one.

            Case 5 - Illness; not work-related

            Case 6 – Visitor to Human Resources; slip, trip, fall, contusion to hip and forehead requiring treatment.

2. Calculate the LWDI .

3. Calculate the total injury/illness incidence rate .

4. Calculate the number-of-lost-workdays rate (5 points).
Answer #1

2. Total work hours for 65 employees - 65 x 2000 = 130,000 work hours per employee per year

Case 1-

a) LWDI - 0 (No workday lost)

b) IR - Incidence Rate - (1 x 200,000)/130,000=1.53

Case 2-

a) LWDI - (5 x 200,000)/130,000 = 7.69

b) IR - Incidence Rate - (1 x 200,000)/130,000 = 1.53